# Refine Neutrino Events Reconstruction with BEiT-3

### v2

@@ I Introduction
IceCube is a neutrino observatory that studies the most extreme phenomena in the universe by observing neutrinos Aartsen, Collaboration et al. (2016). Examples include Gamma-ray Bursts, Active Galactic Nuclei, and recently discovered neutrinos emitted by the Milky Way Collaboration*† et al. (2023). Neutrino Events Reconstruction such as direction reconstruction and energy reconstruction are crucial in the observation of neutrinos. Traditionally, to perform the reconstruction, we would use likelihood methods or Machine Learning algorithm that mostly based on Convolutional Neural Network (CNN) Aiello et al. (2020); Abbasi et al. (2021); Eller et al. (2023), or Graph Neural Network (GNN), Abbasi et al. (2022); Søgaard et al. (2022).
@@
In order to encourage the development of innovative solutions to improve the accuracy and efficiency of Neutrino Event Reconstruction, the Kaggle competition “IceCube – Neutrinos in Deep Ice” Eller (2023) was held between January 19, 2023 and April 19, 2023. Most top solutions use Transformer or GraphNeT integrated with Transformer. Transformer is traditionally used for Natural Language Processing Vaswani et al. (2017), and then Computer Vision Dosovitskiy et al. (2020) by introducing the idea of class token (cls token) and only using the encoder part of the full Transformer. Transformer for vision is typically called Vision Transformer (ViT). Transformer is powerful that it can learn the position information of a sentence or a picture.
@@
In particular, the $2^{\textrm{nd}}$ place solution IceMix uses BEiT v2 Peng et al. (2022) block as encoder. BEiT v2 uses semantic-rich visual tokenizer as the reconstruction target for masked prediction in Computer Vision. However, current models are heavy and the benefits of each other is hard to be gathered.
@@
BEiT-3 Wang et al. (2022a) is a general-purpose multimodal foundation model based on Magneto Wang et al. (2022b). And it can be used by importing TorchScale Ma et al. (2022), which is an open-source toolkit that enables scaling Transformers both efficiently and effectively. ISeeCube is a pure Transformer model based on TorchScale. A pure Transformer struture has been proven powerful in many Machine Learning missions, and in Neutrino Events Reconstruction through this competition. And by utilizing the highly integrated TorchScale, it is easier for refining, maintaining and testing.
@@
For upcoming neutrino observatories such as TRIDENT Ye et al. (2023) and HUNT Huang et al. (2023), considering that (a) they also observe astrophysical neutrinos and (b) their structure is close to IceCube that they are composed of several strings of Digital Optical Modules (DOMs), there’s a great chance that ISeeCube can also be used for these observatories.
@@
The paper is structured as follows: Section II summarize Neutrino Events Reconstruction methods based on likelihood principles, CNN and GNN, and some of the top solutions in this Kaggle competition. We present ISeeCube as a pure Transformer structure for Neutrino Events Reconstruction in section III, and compare regression models and classification models. Followed by section IV we describe the dataset and the training process, and we evaluate the azimuthal and zenithal distribution of the prediction made by regression and classification models. In section V we discuss the potential of ISeeCube and challenges in other missions such as energy reconstruction and cascade/track event classification.
II Related Work
In section II.1 we summarize some of the likelihood methods and Machine Learning methods used in neutrino reconstruction, followed by section II.2 we summarize some of the top solutions in the Kaggle competition.
II.1 Previous Neutrino Events Reconstruction Methods
Previously, likelihood method is widely used for Neutrino Events Reconstruction:
  • SkyLLH Bellenghi, Karl, and Wolf (2023) is a Python-based tool that uses log-likelihood functions.
  • Based on Jammy_flows, by using the differential entropy and the KL-divergence to its maximum entropy approximation Glüsenkamp (2023), the results for showers and muons are interpreted.
Also, there are Machine Learning models based on CNNs or GNNs, for example:
  • CNN model combined with hexagonally shaped kernels Abbasi et al. (2021), which takes advantage of the geometric symmetry in the shape of IceCube.
  • A likelihood-free technique Eller et al. (2023) in which only simulated data is required as input without an explicit assumption of the likelihood function.
  • Deep CNN method Aiello et al. (2020) that treat event as an image in the Neutrino Events Reconstruction for KM3NeT/ORCA.
  • GNN method Abbasi et al. (2022) that treat event as point cloud. It is capable of distinguishing neutrino events from cosmic-ray backgrounds, classify different neutrino event types, and reconstructing the deposited energy, direction and interaction vertex.
  • GraphNeT Søgaard et al. (2022), an open-source python package for Neutrino Events Reconstruction based on GNN. One of the advantages of GraphNeT is that it can be implemented on the original model smoothly, since the GNN process can be defined that it does not change the shape of input.
  • A hybrid approach Hünnefeld (2021) that utilizes generative CNN to approximate the likelihood, which may then be used in a traditional maximum-likelihood setting.
  • CNN model Peterson, Rodriguez, and Hanson (2023) that exploits time and depth translational symmetry in IceCube DeepCore that distinguishing muon neutrinos from other flavors.
Neutrinos from the Milky Way is detected Collaboration*† et al. (2023) by first using CNN Abbasi et al. (2021) to classify cascade events from all the events, and then using hybrid reconstruction method combined with a maximum likelihood estimation. The hybrid reconstruction method uses Neural Network (NN) including symmetries and detector-specific domain knowledge, which is described previously on Hünnefeld (2021), and further discussed in the supplementary materials of Collaboration*† et al. (2023).
II.2 In the Kaggle Competition
@@
In private leaderboard of the Kaggle competition  [^1] https://www.kaggle.com/competitions/icecube-neutrinos-in-deep-ice/leaderboard, these solutions are representative and quite interesting.
@@
  • $1^{\textrm{st}}$: It’s an combination of EdgeConv and Transformer. The Transformer part only has 3-4 layers. The number of total parameters is only 6M, which is the smallest top solutions.
@@
  • $2^{\textrm{nd}}$: The Transformer part is based on BEiT v2 Peng et al. (2022) block. Also, it uses class token and introduces relative spacetime interval bias as Relative Positional Embedding (RPE) Shaw, Uszkoreit, and Vaswani (2018). This solution shows that, with right set of parameters for RPE, embedding dimensions and number of layers, a pure transformer structure works as well. RPE is close to the mechanism that is used in GNN to capture the relationship of nodes in a sequence.
@@
  • $3^{\textrm{rd}}$: It’s a pure Transformer structure, which has two classification heads (for azimuth and zenith) with 128 bins each. Since it’s a classification mission, a Cross Entropy Loss is better than Von Mises-Fisher Loss Kumar and Tsvetkov (2018).
@@
  • $5^{\textrm{th}}$: The GraphNeT part uses LayerNorm as the activation function instead of BatchNorm1d. The Transformer part has two classification heads, each containing $256$ bins for azimuth and zenith. The Transformer consists of $8$ layers, with a model size of $512$ and $8$ attention heads. Additionally, a $3$-layer Bi-directional LSTM model selector is employed to choose between GraphNeT and Transformer. Combining multiple individual models to solve a given problem is called “ensemble” method.
@@
Besides GNN, which wins the Early Sharing Prize and is provided as an example in this competition, some solutions also use Recurrent Neural Network (RNN), which takes into account of time information. Long Short-Term Memory (LSTM) is designed to deal with long-range dependencies. Bidirectional Gate Recurrent Unit (BGRU) is designed to deal with the same problem, with better performance. But as Vaswani et al. (2017) stated, the attention mechanism can be used to solve the same problem and reach better score.
Table 1: Different Machine Learning methods used in some of the Kaggle solutions.
@@
Table 1 summarizes some of the top solutions in the Kaggle competition Bukhari et al. (2023). $\checkmark$” means the architecture is used in the solution, and a specific name represents the model of the architecture. “Ensemble” means whether the final result is a ensemble of several different models by some weighed mechanism.
In summary, Transformer architecture can work really well for having the ability to capture the relationship of both nodes Joshi (2020) and time sequence. What’s more, since the total number of nodes is smaller than vocabulary of a language, ViT is better than Transformer that an event is more close to a picture than sentence.
As for loss function, Von Mises-Fisher Loss Kumar and Tsvetkov (2018) has the advantage that the first Bessel function is the solution for partial differential equations in a cylinder. And IceCube is constructed by several strings, meaning that all the detectors on the same string has the same $x$ and $y$ coordinate, which makes its shape more close to a cylinder. This loss function finds the internal properties of IceCube. However, some solutions indicate that Cross Entropy Loss is better than Von Mises-Fisher Loss. A possible explanation is that these solutions divide the entire sky into several bins and then predict which bin is the most likely, thus it can work really well with Cross Entropy Loss, which is normally used for classification.
III ISeeCube
@@
Figure 1: The structure of ISeeCube. (a) shows how raw data is embedded then put into the encoder, the class token is taken out and transformed into azimuth and zenith as the result, while the register tokens and embedded data are removed. “n” in “emb n” corresponds to the number of pulses in an event, while “n” in “reg n” is a rather arbitrary number. (b) shows the detail of each layer in the Transformer encoder, which is originally figure 2 (a) of Wang et al. (2022b).
  • In section III.1, we summarize previous data representation of neutrino events used in CNNs and GNNs, and conclude that Embedding used in Transformer is a intuitive derivation of the previous methods.
  • In section III.2 we show that by using TorchScale, the lines of code drop sharply. We show that ViT-like model is capable of learning the position and time information of neutrino events, and future new models or upgrades on Transformer Encoder and ViT has the potential to be used on Neutrino Events Reconstruction. Simple methods work really well, such as Sinusoidal Position Embedding, Relative Position Embedding (RPE) Raffel et al. (2020), Class Token and Register Tokens Darcet et al. (2023).
  • In section III.3 we compare regression and classification models by using different Multilayer Perceptron (MLP) on the class token.
  • In section III.4 we compare Mean Squared Error (MSE) Loss and Cross Entropy (CE) Loss used in regression and classification models, respectively. We also discuss the usage of the official metric in the Kaggle competition.
III.1 Embedding: Data Representation
@@
  • Neutrino event is loosely arranged, meaning that (a) the distance between each DOM is not the same, and (b) not all the DOMs detect photons in an event, see figure 2
@@
CNNs: In Peterson, Rodriguez, and Hanson (2023) and Abratenko et al. (2021), neutrino event is represented as $\mathbb{R}^{S\times N\times T}$, where $S$ is the number of strings, $N$ is the number of DOMs on each string and $T$ is time segments. Time segments is the total time divided by time periods. And the number in this matrix represents charge. This matrix is then put into a CNN. For example, in Abratenko et al. (2021) the CNN is based on ResNet He et al. (2016).
@@
Transformers: Same as GNN method, the neutrino event is also represented as $\mathbb{R}^{N\times P}$. Firstly, when the total number of pulses is bigger than the given $N$, we would choose auxiliary False first, then True; when the total number of pulses is smaller than $N$, the empty space will be filled with the position of the first DOM (the choice of the first DOM is rather arbitrary) and time, charge, auxiliary would be zero. For more information on the dataset and how to normalize the event, see section IV.1. Then Sinusoidal Position Embedding is used. In Vaswani et al. (2017), Sinusoidal Position Embedding assigns a unique fixed vector to each position in the input sequence. These vectors are computed using sine and cosine functions with different frequencies. In neutrino events it is introduced for the same purpose, projecting $\mathbb{R}^{N\times P}$ to a higher dimension. In our case, for example, one of the properties $x$ (shape $\mathbb{R}^{N\times 1}$), is projected to $\mathbb{R}^{N\times N}$ by:
$PE_{(i,j)}$ $=\sin\left(\frac{pos(i)}{10000^{j/N}}\right),j=0,1,\cdots,\frac{N}{2}-1$ (1)
@@
where $pos(i)$ is each element in $\mathbb{R}^{N\times 1}$. Then the rest of the properties will go through the same process and all the embedded elements will be concatenated together ($x$, $y$, $z$ will be concatenated first, then the rest of the properties). Then we use a MLP to produce the final embedded event as $\mathbb{R}^{N\times d_{model}}$. This is the input of the Transformer Encoder.
@@ III.2 Transformer Encoder
  • Out-of-the-box: An encoder structure can be used by few lines of code. A lot of methods can be tested by adjusting the config. For example, Relative Position Embedding (RPE). While in the $2^{\textrm{nd}}$ place solution, the layers for encoder have be rewritten.
@@
  • Only relay on few dependencies including PyTorch, timm and fairscale, while GNNs rely on PyTorch, PyTorch Cluster, PyTorch Scatter, PyTorch Sparse and
    PyTorch Geometric.
@@
  • In Feed Forward Network (FFN), A sub_LayerNorm (sub_LN) is added to stabilize models with deeper layers. Also, the default activation function in FFN is now GELUs Hendrycks and Gimpel (2016) rather than ReLUs, which combines dropout zoneout and ReLUs to alleviate the vanishing gradient problem. Also, a dropout layer is added to alleviate the same problem.
@@
Relative Position Embedding (RPE) Raffel et al. (2020) is presented as a new method to learn the relative position information. In the $2^{\textrm{nd}}$ place solution, RPE is added by introducing Special Relativity bias and writing new layers and blocks with this bias. In TorchScale, RPE can be added easily by adjusting few parameters, though without Special Relativity bias.
@@
In ViT Dosovitskiy et al. (2020), Class Token is concatenated to the embedding of the picture to learn the position information. In ISeeCube, Class Token is concatenated to the embedding of a neutrino event for the same purpose, to learn the position information of neutrino events.
@@
Register Tokens Darcet et al. (2023) are used for smoother feature maps and attention maps for downstream visual processing when finetuning ViT models. In ISeeCube, Register Tokens are concatenated to the embedding of a neutrino event for the same purpose. Also, we notice that when the number of class token is $1$, the number of register tokens is $3$, and we pick $196$ pulses in an event, the total sequence length been put into the encoder is $200$. The “warp size” of a CUDA core is the number of threads in a warp executed simultaneously on a CUDA core, which is normally $32$. When multiple threads within a warp access adjacent memory locations, the GPU can coalesce these memory transactions into a single memory transaction, resulting in improved memory throughput. $200$ is a multiple of $8$, which is more CUDA-friendly than the case with only class token, $197$, which is a prime number.
III.3 Multilayer Perceptron: Regression and Classification

### v4

@@ 1 Introduction
IceCube is a neutrino observatory that studies the most extreme phenomena in the universe by observing neutrinos [1]. Examples include Gamma-ray Bursts, Active Galactic Nuclei, and recently discovered neutrinos emitted by the Milky Way [2]. Neutrino Events Reconstruction such as direction reconstruction and energy reconstruction are crucial in the observation of neutrinos. Traditionally, to perform the reconstruction, we would use maximum likelihood methods or Machine Learning models that mostly based on Convolutional Neural Networks (CNNs) [3, 4, 5, 6, 7], or Graph Neural Networks (GNNs), [8, 9].
@@
In order to encourage the development of innovative solutions to improve the accuracy and efficiency of Neutrino Events Reconstruction, the Kaggle competition “IceCube – Neutrinos in Deep Ice” [10] was held between January 19, 2023 and April 19, 2023. Most top solutions use Transformer or GraphNeT integrated with Transformer. Transformer is traditionally used for Natural Language Processing [11], and then Computer Vision [12] by introducing the idea of class token (cls token) and only using the encoder part of the full Transformer, which is called Vision Transformer (ViT). Transformer is powerful that it can learn the position information of a sentence or a picture.
@@
In particular, the $2^{\textrm{nd}}$ place solution IceMix uses BEiT v2 [13] block as encoder. BEiT v2 uses semantic-rich visual tokenizer as the reconstruction target for masked prediction in Computer Vision. However, current models are heavy and the benefits of each other is hard to be gathered.
@@
BEiT-3 [14] is a general-purpose multimodal foundation model based on Magneto [15]. And it can be used by importing TorchScale [16], which is an open-source toolkit that enables scaling Transformers both efficiently and effectively. ISeeCube is a pure Transformer model based on TorchScale. A pure Transformer structure has been proven powerful in many Machine Learning missions, and in Neutrino Events Reconstruction through this competition. And by utilizing the highly integrated TorchScale, it is easier for refining, maintaining and testing.
@@
For upcoming neutrino observatories such as TRIDENT [17] and HUNT [18], considering that (a) they also observe astrophysical neutrinos and (b) their structure is close to IceCube that they are also composed of several strings of Digital Optical Modules (DOMs), there’s a great chance that ISeeCube can also be used for these observatories.
@@
The paper is structured as follows: Section 2 summarize Neutrino Events Reconstruction methods and some of the top solutions in this Kaggle competition. We present ISeeCube as a pure Transformer structure for Neutrino Events Reconstruction in section 3, and compare regression models and classification models. Followed by section 4 we describe the dataset and the training process, and we evaluate the azimuthal and zenithal distribution of the prediction made by regression and classification models. In section 5 we discuss the potential of ISeeCube and challenges in other missions such as energy reconstruction and cascade/track event classification.
2 Related Work
Previously, likelihood method [19, 20] is widely used for Neutrino Events Reconstruction. Also, there are Machine Learning models based on CNNs [3, 4, 5, 6, 7] or GNNs [8, 9]. These two Neural Networks (NNs) represent two different approaches of data representation, see section 3.1. Notably, neutrinos from the Milky Way are detected [2] by first using CNN [3] to classify cascade events from all the events, and then using hybrid reconstruction method combined with a maximum likelihood estimation [6].
@@ 2 Related Work
In private leaderboard of the Kaggle competition  [^2] https://www.kaggle.com/competitions/icecube-neutrinos-in-deep-ice/leaderboard, these top solutions are representative:
@@
  • $1^{\textrm{st}}$: It contains 3-4 layers and the number of total parameters is only 6M (one of the smallest top solutions). Each of these layer is composed of an EdgeConv layer and a Transformer Encoder layer sequentially.
@@
  • $2^{\textrm{nd}}$: The Transformer part is based on BEiT v2 [13] block. Also, it uses class token and introduces relative spacetime interval bias as Relative Positional Embedding (RPE) [21]. It has two Encoders, the first Encoder uses RPE and the second does not. This solution shows that, with right set of parameters for RPE, embedding dimensions and number of layers, a pure Transformer structure works as well. RPE is close to the mechanism that is used in GNNs to capture the relationship of nodes in a sequence. This solution also uses GNNs as the first layer of the model to further improve the score.
@@
  • $3^{\textrm{rd}}$: It’s a pure Transformer structure, which has two classification heads (for azimuth and zenith) with 128 bins each. Since it’s a classification mission, a Cross Entropy Loss is better than Von Mises-Fisher Loss [22].
@@
  • $5^{\textrm{th}}$: The GraphNeT part uses LayerNorm as the activation function instead of BatchNorm1d. The Transformer part has two classification heads, each containing $256$ bins for azimuth and zenith. The Transformer consists of $8$ layers, with a model size of $512$ and $8$ attention heads. Additionally, a $3$-layer Bidirectional LSTM model selector is employed to choose between GraphNeT and Transformer. Combining multiple individual models to solve a given problem is called “ensemble” method.
@@
Besides GNNs, which wins the Early Sharing Prize and is provided as an example in this competition, some solutions also use Recurrent Neural Networks (RNNs), which take into account of time information and treat neutrino events as time sequences. Long Short-Term Memory (LSTM) and Gate Recurrent Unit (GRU) are designed to deal with long-range dependencies. But as the original Transformer paper [11] stated, a pure Transformer model instead of RNNs can be used to solve the same problem and reach better score.
Table 1: Machine Learning methods in some of the Kaggle solutions.
@@
Table 1 summarizes some of the top solutions in the Kaggle competition [23]. $\checkmark$” means the architecture is used in the solution, and a specific name represents the specific model. “Ensemble” means whether the final result is a ensemble of several different models by some weighed mechanism.
In summary, Transformer architecture can work really well for having the ability to capture the information of both graphs [24] and time sequences. What’s more, since the total number of nodes in a graph is smaller than vocabulary of a language, ViT is better than Transformer that an event is closer to a picture than sentence.
As for loss function, Von Mises-Fisher Loss [22] has the advantage that the first Bessel function is the solution for partial differential equations in a cylinder. And IceCube is constructed by several strings, meaning that all the detectors on the same string has the same $x$ and $y$ coordinate, which makes its shape closer to a cylinder. This loss function finds the internal properties of IceCube. However, some solutions indicate that Cross Entropy Loss is better than Von Mises-Fisher Loss. A possible explanation is that these solutions divide the azimuth and zenith into several bins and then predict which bin is the most likely, thus it can work really well with Cross Entropy Loss, which is normally used for classification missions.
3 ISeeCube
We noticed that in the $2^{\textrm{nd}}$ place solution, new layers and blocks have to be re-written to apply Special Relativity bias and the structure of the model is not clear (see section 3.2). Also, MSE Loss and CE Loss are simpler than VMF Loss (see section 3.4). In order to enhance structural clarity and simplicity, reduce lines of code, and investigate the feasibility of implementing Transformers and their upgrades for Neutrino Events Reconstruction, we made several enhancements to the $2^{\textrm{nd}}$ place solution, see table 2. The code for embedding (see section 3.1) and data-loading (see section 4.1) are inherited from the $2^{\textrm{nd}}$ place solution. The rest of the code are re-written.
@@ 3 ISeeCube
Figure 1: The structure of ISeeCube. (a) shows how raw data is embedded then put into the encoder, the class token is taken out and transformed into azimuth and zenith as the result, while the register tokens and embedded data are removed. “n” in “emb n” corresponds to the number of pulses in an event, while “n” in “reg n” is a rather arbitrary number. (b) shows the detail of each layer in the Transformer Encoder (originally figure 2 (a) in Magneto [15]).
Table 2: Differences and commonalities between the $2^{\textrm{nd}}$ place solution and ISeeCube.
$2^{\textrm{nd}}$ place solution ISeeCube
Embedding Sinusoidal Positional Embedding
Transformer Encoder BEiT v2 BEiT-3 (TorchScale)
Number of Encoders 2 1 or 2
RPE $\checkmark$, Special Relativity Bias $\checkmark$
Token Class Token Class Token, Register Tokens
Loss VMF MSE or CE
Data-Loading Chunk-Based Loading, Batch-Based Loading
  • In section 3.1, we summarize previous data representation of neutrino events for CNNs and GNNs, and conclude that Embedding in Transformer is a intuitive derivation of the previous methods.
  • In section 3.2 we show that by using TorchScale, the lines of code drop sharply. We show that ViT-like model is capable of learning the position and time information of neutrino events, and future new models or upgrades on Transformer Encoder and ViT has the potential to be used on Neutrino Events Reconstruction. Simple methods work really well, such as Sinusoidal Positional Embedding, Relative Position Embedding (RPE) [25], Class Token and Register Tokens [26].
  • In section 3.3 we compare regression and classification models by using different Multilayer Perceptrons (MLPs) on the class token.
  • In section 3.4 we compare Mean Squared Error (MSE) Loss and Cross Entropy (CE) Loss used in regression and classification models, respectively. We also discuss the usage of the Kaggle competition metric.
3.1 Embedding: Data Representation
@@
  • Neutrino event is loosely arranged, meaning that (a) the distance between each DOM is not the same, and (b) not all the DOMs detect photons in an event, see figure 2.
@@
CNNs: In some CNNs models [7, 27], neutrino event is represented as $\mathbb{R}^{S\times N\times T}$, where $S$ is the number of strings, $N$ is the number of DOMs on each string and $T$ is time segments. Time segments is the total time divided by time periods. And the value in this tensor represents charge. This tensor is then put into a CNN. For example, in [27] the CNN is based on ResNet [28].
@@
Transformers: Same as the GNN method, the neutrino event is also represented as $\mathbb{R}^{N\times P}$. Firstly, when the total number of pulses is bigger than the given $N$, we would choose auxiliary False first, then True; when the total number of pulses is smaller than $N$, the empty space will be filled with the position of the first DOM (the choice of the first DOM is rather arbitrary) and time, charge, auxiliary would be zero. For more information on the dataset and how to normalize the event, see section 4.1. Then Sinusoidal Positional Embedding is used. In the original Transformer [11], Sinusoidal Positional Embedding assigns a unique fixed vector to each position in the input sequence. These vectors are computed using sine and cosine functions with different frequencies. In neutrino events it is introduced for the same purpose, projecting $\mathbb{R}^{N\times P}$ to a higher dimension. In our case, for example, one of the properties $x$ (shape $\mathbb{R}^{N\times 1}$), is projected to $\mathbb{R}^{N\times N}$ by:
$PE_{(i,j)}$ $=\sin\left(\frac{pos(i)}{10000^{j/N}}\right),j=0,1,\cdots,\frac{N}{2}-1$ (3.1)
@@
where $pos(i)$ is each element in $x$. Then the rest of the properties will go through the same process and all the embedded elements will be concatenated together ($x$, $y$, $z$ will be concatenated first, then the rest of the properties). Then we use a MLP to produce the final embedded event as $\mathbb{R}^{N\times d_{model}}$. This is the input of the Transformer Encoder.
@@ 3.2 Transformer Encoder
  • Out-of-the-box: An encoder structure can be imported by few lines of code. A lot of methods can be tested by adjusting the config. For example, Relative Position Embedding (RPE). While in the $2^{\textrm{nd}}$ place solution, the layers for encoder have be rewritten. Specifically, besides class Block from BEiT v2, several new classes Attention_rel, Block_rel, Rel_ds and LocalBlock have to be modified or introduced to apply Special Relativity bias.  [^3] Please refer to https://github.com/DrHB/icecube-2nd-place/blob/main/src/models.py and https://github.com/microsoft/unilm/blob/master/beit2/modeling_finetune.py.
@@
  • Only rely on few dependencies including PyTorch, timm and fairscale, while GNNs rely on PyTorch, PyTorch Cluster, PyTorch Scatter, PyTorch Sparse and PyTorch Geometric.
@@
  • In Feed Forward Network (FFN), A sub_LayerNorm (sub_LN) is added to stabilize models with deeper layers. The default activation function in FFN is now GELUs [29] rather than ReLUs, which combines dropout zoneout and ReLUs to alleviate the vanishing gradient problem. And a dropout layer is added to alleviate the same problem.
@@
Relative Position Embedding (RPE) [25] is presented as a new method to learn the relative position information. In the $2^{\textrm{nd}}$ place solution, RPE is added by introducing Special Relativity bias and writing new layers and blocks with this bias. In TorchScale, RPE can be added easily by adjusting few parameters, though without Special Relativity bias.
@@
In ViT [12], Class Token is concatenated to the embedding of the picture to learn the position information. In ISeeCube, Class Token is concatenated to the embedding of a neutrino event for the same purpose, to learn the position information of neutrino events.
@@
Register Tokens [26] are used for smoother feature maps and attention maps for downstream visual processing when finetuning ViT models. In ISeeCube, Register Tokens are concatenated to the embedding of a neutrino event for the same purpose. For image classification, the predicted class corresponds to the index with the largest probability; for Neutrino Events Reconstruction, the model directly predict the value itself. Thus smoother feature maps and attention maps are really helpful. Also, in terms of memory usage, we notice that when the number of class token is $1$, the number of register tokens is $3$, and we pick $196$ pulses in an event, the total sequence length been put into the encoder is $200$. The “warp size” of a CUDA core is the number of threads in a warp executed simultaneously on a CUDA core, which is normally $32$. When multiple threads within a warp access adjacent memory locations, the GPU can coalesce these memory transactions into a single memory transaction, resulting in improved memory throughput. $200$ is a multiple of $8$, which is more CUDA-friendly than the case with only class token, $197$, which is a prime number.
3.3 Multilayer Perceptron: Regression and Classification